OpenStudy (anonymous):

multiply the following two binomials together (t+3) and (2t-4)

OpenStudy (anonymous):

(t+3)(2t-4)=t(2t-4)+3(2t-4)=2t^2-4t+6t-12=2t^2+2t-12

OpenStudy (anonymous):

Try splitting the first binomial and multiplying each of its terms in turn by the second... (t+3)(2t-4) = t(2t-4) +3(2t-4) = 2t^2 - 4t + 6t - 12 Now gather like terms and simplify... 2t^2 + 2t - 12